OpenStudy (anonymous):

Match each equation with the description of the first step you can take to solve the equation.

OpenStudy (anonymous):

K/6 - 8/15 = -4/15 A.) ADD 4/15 TO BOTH SIDES -4/15 + K/6 = -8/15 B.) SUBTRACT 4/15 FROM BOTH SIDES K/6 + 4/15 = -8/15 C.) SUBTRACT 8/15 FROM BOTH SIDES 8/15 + K/6 += -4/15 D.) ADD 8/15 TO BOTH SIDES

OpenStudy (anonymous):

1. D 2. A 3. B 4. C

OpenStudy (anonymous):

K/6 - 8/15 = -4/15 (add 8/15 to both sides) D -4/15 + K/6 = -8/15 (add 4/15 to both sides) A K/6 + 4/15 = -8/15 (subtract 4/15) B 8/15 + K/6 = -4/15 (subtract 8/15 ) C I agree